# Variational Pseudo Marginal Methods for Jet Reconstruction in Particle Physics

### v1

@@ 4 Experiments
The industry standard in particle physics uses agglomerative clustering techniques, which are greedy [Cacciari et al., 2012]. Beam Search provides a straightforward and significant improvement. Thus, we consider both Greedy and Beam Search as standard, relevant and efficient baselines, also applied in cited works Greenberg et al. [2020, 2021]. We simulated 100 jets using Ginkgo running comparisons with Greedy Search, Beam Search and Cluster Trellis.
MAP estimate.
@@
Fig. 5 (left) provides a scatterplot comparing log-conditional likelihood values. Across 100 simulated jets, Vncsmc with $K,M=(256,1)$ returns a higher likelihood on all 100 cases against Greedy Search (left) and 99 cases against Beam Search (right). Notably, Vncsmc simultaneously conducts inference and $\lambda$ learning, a feature lacking in Greedy Search and Beam Search, which rely on the user providing $\lambda$ values. Furthermore, Vncsmc yields probability distributions over topologies, while Greedy Search and Beam Search yield single topologies. Fig. 5 (right) shows the log conditional likelihood values $\hat{p}(X|\tau,\lambda)$ for Vcsmc (blue) and Vncsmc (red) each with $K=\{8,256\}$ (and $M=1$ for Vncsmc) samples averaged across 5 random seeds. Greater values of $K$ result in higher values, while reducing stochastic gradient noise. Vcsmc with $K\geq 8$ explores higher probability spaces than the simulator, as depicted by the dotted green reference trace. Vncsmc achieves convergence in fewer epochs than Vcsmc and yields higher log-conditional likelihood values, all while maintaining lower stochastic gradient noise. Notably, even Vncsmc with $(K,M)=(8,1)$ in the top-left plot (red) outperforms Vcsmc with $K=256$ in the bottom-right plot (blue). Next, we compare in Fig. 6 (left) the log conditional likelihood returned by VCSMC, Greedy Search and Beam Search with the exact MAP clustering value calculated with the cluster trellis technique described in  Greenberg et al. [2020], for a dataset of 100 jets. We see that VCSMC returns high quality hierarchies, with log likelihood values close to the exact ones.
Running Time and Complexity.
@@
We generated jets with $N=\{4,\cdots,64\}$ leaf nodes and profiled the running time of Vcsmc, Cluster Trellis, Greedy Search and Beam Search averaged across 3 random seeds. All experiments were performed on a Google Cloud Platform n1-standard-4 instance with an Intel Xeon CPU 4 vCPUs and 15 GB RAM without leveraging GPU utilization. Fig. 6 (left center) reports the running times on a log scale in seconds. Vcsmc is an order of magnitude faster than Beam Search on $N=20$ leaf nodes. Beam search entails managing a list of log-likelihood pairs at each level and for each beam size $b$. The given list is sorted, iterated over, and selectively only a single topology is retained at a time. This incurs a complexity of $\mathcal{O}(b^{2}N\log b+bN^{3}\log N)$. Typically $b>N$, in which case the complexity can be prohibitively slow, but for $b\log<N^{2}\log N$, the complexity becomes $\mathcal{O}(bN^{3}\log N)$. In contrast, Ncsmc is $\mathcal{O}(KN^{3}M)$ and the Csmc is $\mathcal{O}(KNM)$, where $K,N$ and $M$ denote the number of Monte Carlo samples, the number of leaf nodes (observed particles), and the number of subsamples.
@@
Fig. 6 (center right) illustrates convergence of the conditional Csmc and Ncsmc likelihood to the cluster trellis marginal likelihood as $K$ increases. Finally, Fig. 6 (right) illustrates the inferred Log-Normal Pseudo-Marginal Distribution for the parameters $\mu=(\mu_{1},\mu_{2})$ of the Heavy Resonance Jet, estimated through Ncsmc. Contours of the log-conditional likelihood are shown with stochastic gradient steps taken on $\mathcal{L}_{\textrm{NCSMC}}$ highlighted in red.
5 Conclusion
We have introduced the first adaptation of Csmc for unbiased and consistent jet reconstruction, proposing approximate posteriors and VI for both point and distribution estimators. Additionally, we’ve established connections between Vcsmc and variational pseudo-marginal methods, showcasing significant improvements in speed and accuracy, laying the groundwork for wider adoption of variational methods in collider data analyses.
Acknowledgements.

### v2

@@ 4 Experiments
The industry standard in particle physics uses agglomerative clustering techniques, which are greedy (Cacciari et al., 2012). Beam Search provides a straightforward and significant improvement. Thus, we consider both Greedy and Beam Search as standard, relevant and efficient baselines, also applied in cited works Greenberg et al. (2020; 2021). We simulated 100 jets using Ginkgo running comparisons with Greedy Search, Beam Search and Cluster Trellis.
4.1 MAP estimate
@@
Fig. 5 (left) provides a scatterplot comparing log-conditional likelihood values. Across 100 simulated jets, Vncsmc with $K,M=(256,1)$ returns a higher likelihood on all 100 cases against Greedy Search (left) and 99 cases against Beam Search (center). Notably, Vncsmc simultaneously conducts inference and $\lambda$ learning, a feature lacking in Greedy Search and Beam Search, which rely on the user providing $\lambda$ values. Furthermore, Vncsmc yields probability distributions over topologies, while Greedy Search and Beam Search yield single topologies.
Fig. 5 (right) shows the log conditional likelihood $\log\hat{p}(X|\tau,\lambda)$ for Vcsmc (blue) and Vncsmc (red) with $K={256}$ (and $M=1$ for Vncsmc) samples averaged across 5 random seeds. Vncsmc converges in fewer epochs and achieves higher likelihood values with lower gradient noise. Additional experiments illustrating the effect of $K$ appear in Fig. 9 in the Appendix. Larger $K$ yields higher $\log\hat{p}(X|\tau,\lambda)$ values and reduces stochastic gradient noise.
Figure 6: Left: Csmc and Ncsmc converge to the exact marginal likelihood as $K$ increases (log scale). Center: Log conditional likelihood returned by Vcsmc compared with the exact MAP clustering returned by the cluster trellis Greenberg et al. (2020), Greedy Search and Beam Search, for a dataset of 100 simulated jets generated with Ginkgo. Even with $K=256$ and $N=12$, Vcsmc closely approximates exact cluster trellis values for specific jets. Right: A comparison of the running times, highlighting that Vcsmc on $N=64$ significantly outperforms cluster trellis on $N=12$. As $N$ increases, the cluster trellis quickly becomes impractical due to its exponential complexity.
4.2 Running Time and Complexity.
@@
We generated jets with $N=\{4,\cdots,64\}$ leaf nodes and profiled the running time of Vcsmc, Cluster Trellis, Greedy Search and Beam Search averaged across 3 random seeds. All experiments were performed on a Google Cloud Platform n1-standard-4 instance with an Intel Xeon CPU 4 vCPUs and 15 GB RAM without leveraging GPU utilization.
@@
Fig. 6 (left) illustrates convergence of the Csmc and Ncsmc conditional likelihood to the exact cluster trellis marginal likelihood as $K$ increases. Fig. 7 plots the inferred Log-Normal Pseudo-Marginal Distribution for the parameters $\mu=(\mu_{1},\mu_{2})$ of the Heavy Resonance Jet, estimated through Ncsmc. Contours of the log-conditional likelihood are shown with stochastic gradient steps taken on $\mathcal{L}_{\textrm{NCSMC}}$ highlighted in red.
Next, we compare in Fig. 6 (center) the log conditional likelihood returned by VCSMC, Greedy Search and Beam Search with the exact Maximum a Posteriori (MAP) clustering value calculated with the cluster trellis technique described in  Greenberg et al. (2020), for a dataset of 100 jets. We see that VCSMC returns high quality hierarchies, with log likelihood values close to the exact ones.
Fig. 6 (right) reports the running times on a log scale in seconds. Vcsmc is an order of magnitude faster than Beam Search on $N=20$ leaf nodes. Beam search entails managing a list of log-likelihood pairs at each level and for each beam size $b$. The given list is sorted, iterated over, and selectively only a single topology is retained at a time. This incurs a complexity of $\mathcal{O}(b^{2}N\log b+bN^{3}\log N)$. Typically $b>N$, in which case the complexity can be prohibitively slow, but for $b<N^{2}\log N$, the complexity becomes $\mathcal{O}(bN^{3}\log N)$. In contrast, Ncsmc is $\mathcal{O}(KN^{3}M)$ and the Csmc is $\mathcal{O}(KNM)$, where $K,N$ and $M$ denote the number of Monte Carlo samples, the number of leaf nodes (observed particles), and the number of subsamples.
Figure 7: Inferred log-normal pseudo-marginal distribution for the parameters $\hat{\mu}=(\hat{\mu}_{1},\hat{\mu}_{2})$ of the heavy resonance jet, estimated using Ncsmc. Contours represent the log-conditional likelihood, with stochastic gradient steps on $\mathcal{L}_{\textrm{NCSMC}}$ highlighted in red.
5 Discussion
In all experiments, $t_{cut}$ is an exogenous variable, and in full physics simulations its value is determined by the energy scale where there is a qualitative change in the theory that describes the process, i.e. we switch from a showering to a hadronization process, described by different models of physics. The difference between the energy of the initial state particle (root of the binary tree) and this energy scale ($t_{cut}$) affects the number of final state particles in the Monte Carlo sampling. Given that Ginkgo does not include a full quantum chromodynamics (QCD) modeling of the splitting likelihood, we choose $t_{cut}$ to control the distribution on the number of final state particles and mimic full physics simulation processes. As such, the exact value is not relevant and the developed algorithms capabilities are independent of it.
Simulators rooted in QCD present significant challenges due to their complex likelihoods. Rewriting these simulators is a substantial endeavor, demanding considerable expertise and effort, with specialists often dedicating entire careers to mastering the intricacies of QCD. Real-world data utilization necessitates strict adherence to the models embedded within these simulators, further complicating the task. Moreover, the lack of currently available powerful quantum computers adds another layer of difficulty, prompting high-energy researchers, such as those at Iris-Hep, to rely on approximations like Ginkgo (Cranmer et al., 2019a).
The Csmc method (Wang et al., 2015) and the Variational Combinatorial Sequential Monte Carlo (Vcsmc) method  (Moretti et al., 2021) both use a specific generative model for biology based on a continuous time Markov chain. Our paper represents the first adaptation of Sequential Monte Carlo methods for jet reconstruction in particle physics. This is achieved by adapting the Ginkgo model and redefining the parent node splitting likelihood recursively (Eq. 12) to ensure that, at the final Csmc rank event, the support of the proposal distribution aligns with that of the target distribution.
5.1 Conclusion
Variational and pseudo-marginal methods are broadly applicable across a range of combinatorial and continuous problems. In Bayesian phylogenetics, these methods facilitate efficient estimation of marginal likelihoods over an equivalent factorial search space of tree topologies, $(2N-3)!!$!, allowing for uncertainty in both tree structure and branch lengths (Moretti et al., 2021; Zhang & Matsen IV, 2019; Zhang et al., 2021). Additionally, in social and biological network analysis, variational techniques are central to hierarchical clustering and stochastic block models for complex, structured latent spaces (Zhou, 2015; Zhang & IV, 2018; Greenberg et al., 2021). Monte Carlo methods, widely used to model self-avoiding random walks on combinatorial structures, provide insight into diffusion and spatial distribution phenomena that bridge into statistical mechanics and more complex particle interaction models (Madras & Slade, 1996; Grosberg et al., 2006). These techniques also extend to probabilistic graphical models and Bayesian networks, where factorized approximations reduce computational burden, supporting inference on high-dimensional data (Zhang & IV, 2018).
We have introduced the first adaptation of Csmc for unbiased and consistent jet reconstruction, proposing approximate posteriors and variational inference (VI) techniques for both point and distribution estimators. Our approach significantly improves both speed and accuracy, paving the way for broader adoption of variational methods in collider data analyses. This work not only provides a robust framework for jet reconstruction but also sets the stage for future advancements in the application of advanced approximate inference methods to high-energy physics experiments.
Acknowledgements
@@
Appendix A Ginkgo Generative Model
The Ginkgo generative process is outlined below:
Algorithm 1 Toy Parton Shower Generator
0:  parent momentum $p^{\mu}_{p}$, parent mass squared $t_{P}$, cut-off mass squared $t_{\text{cut}}$, rate for the exponential distribution $\lambda$, binary tree $tree$
1:  Function NodeProcessing($p^{\mu}_{p}$, $t_{P}$, $t_{\text{cut}}$, $\lambda$, $tree$)
2:  Add parent node to $tree$.
3:  if $t_{P}>t_{\text{cut}}$ then
4:     Sample $t_{L}$ and $t_{R}$ from the decaying exponential distribution.
5:     Sample a unit vector from a uniform distribution over the 2-sphere.
6:     Compute the 2-body decay of the parent node in the parent rest frame.
7:     Apply a Lorentz boost to the lab frame to each child.
8:     call NodeProcessing($p^{\mu}_{p}$, $t_{L}$, $t_{\text{cut}}$, $\lambda$, $tree$)
9:     call NodeProcessing($p^{\mu}_{p}$, $t_{R}$, $t_{\text{cut}}$, $\lambda$, $tree$)
10:  end if
The 2-body decay in the parent rest frame is defined using momentum $p_{p}^{\mu}=p_{L}^{\mu}+p_{R}^{\mu}=(\sqrt{s},0,0,0)$. Due to energy-momentum conservation the child energies are given by
$E_{L}=\frac{\sqrt{s}}{2}\left(1+\frac{t_{L}}{s}-\frac{t_{R}}{s}\right)$ (18)
$E_{R}=\frac{\sqrt{s}}{2}\left(1+\frac{t_{R}}{s}-\frac{t_{L}}{s}\right)$ (19)
and the magnitude of their 3-momentum is defined
$|\vec{p}|=\frac{\sqrt{s}}{2}\bar{\beta}=\frac{\sqrt{s}}{2}\sqrt{1-\frac{2(t_{L}+t_{R})}{s}+\frac{(t_{L}-t_{R})^{2}}{s^{2}}}$ (20)
The left and right child momentum are given by $p_{L}^{\mu}=(E_{L},\vec{p})$ and $p_{R}^{\mu}=(E_{R},-\vec{p})$ in the parent rest frame. The Lorentz boost $\gamma=\frac{E_{p}}{\sqrt{t_{p}}}$ and $\gamma\beta=|\vec{p}_{p}|/\sqrt{t_{p}}$. For more information see Cranmer, Kyle et al. (2021); Cranmer et al. (2019b).
Appendix B Combinatorial Sequential Monte Carlo
We provide an overview of the Csmc algorithm from Wang et al. (2015).
B.1 Partial States and the Natural Forest Extension
Definition 1 (Partial State).
A rank $r\in\{0,\cdots,N-1\}$ partial state, symbolized as $s_{r}={(t_{i},X_{i})}$, represents a collection of rooted trees and adheres to the following three conditions:
  1. 1.
    Partial states of different ranks are disjoint, meaning that for any two distinct ranks, $r$ and $s$, there is no overlap between the sets of partial states, written as $\forall r\neq s$, $\mathcal{S}_{r}\cap\mathcal{S}_{s}=\emptyset$.
  2. 2.
    The set of partial states at the smallest rank consists of only one element, denoted as $S_{0}={\bot}$.
  3. 3.
    The set of partial states at the final rank $R=N-1$ corresponds to the target space $\mathcal{X}$.
The likelihood, as represented in Eq. 12, and the probability measure $\pi$ are specifically defined within the scope of the target space, denoted as $\mathcal{S}_{R}=\mathcal{X}$. It’s important to note that these definitions apply exclusively to the target space of trees and not to the broader sample space encompassing partial states, denoted as $\mathcal{S}_{r<R}$, which consists of forests containing disjoint trees. The Sum-Product algorithm is primarily utilized to derive a maximum likelihood estimate for a tree. However, partial states are explicitly characterized as collections of these disjoint trees or leaf nodes. To extend the target measure $\pi$ to encompass the sample space $\mathcal{S}_{r<R}$, a practical approach is to treat all elements of the jump chain as trees, as elaborated in Wang et al. (2015).
Definition 2 (Natural Forest Extension).
The natural forest extension, denoted as Nfe, expands the target measure $\pi$ into forests by forming a product over the trees contained within the forest:
$\pi(s)\coloneqq\prod\limits_{(t_{i},X_{i})}\pi_{Y_{i}(x_{i})}(t_{i})\,.$ (21)
One notable advantage of the Nfe is its ability to transmit information from non-coalescing elements to the local weight update.
Algorithm 2 Combinatorial Sequential Monte Carlo
  Input: $\mathbf{Y}=\{Y_{1},\cdots,Y_{M}\}\in\Omega^{NxM}$, $\theta$
1:  Initialization. $\forall k$, $s_{0}^{k}\leftarrow\perp$, $w_{0}^{k}\leftarrow 1/K$.
2:  for $r=0$ to $R=N-1$ do
3:     for $k=1$ to $K$ do
4:        Resample
$\mathbb{P}(a_{r-1}^{k}=i)=\frac{w_{r-1}^{i}}{\sum_{l=1}^{K}w_{r-1}^{l}}$
5:        Extend partial state
$s_{r}^{k}\sim q(\cdot|s_{r-1}^{a_{r-1}^{k}})$
6:        Compute weights
$w_{r}^{k}=w(s_{r-1}^{a_{r-1}^{k}},s_{r}^{k})=\frac{\pi(s_{r}^{k})}{\pi(s_{r-1}^{a_{r-1}^{k}})}\cdot\frac{\nu^{-}(s_{r-1}^{a_{r-1}^{k}})}{q(s_{r}^{k}|s_{r-1}^{a_{r-1}^{k}})}$
7:     end for
8:  end for
9:  Output: $s_{R}^{1:K}$ , $w_{1:R}^{1:K}$
Appendix C Nested Combinatorial Sequential Monte Carlo
We provide a review of the Ncsmc algorithm from Moretti et al. (2021). The Ncsmc method performs a standard Resample step (line 4), similar to Csmc methods, iterating over rank events. In each iteration, Ncsmc explores all possible one-step ahead topologies ((${N-r\choose 2}$) and samples sub-branch lengths for each of them (line 5-7). Importance sub-weights or potential functions are evaluated for these sampled look-ahead states (line 8). The ancestral partial state is then extended to a new partial state by choosing a topology and branch length based on their respective weights (line 11). The final weight for each sample is calculated by averaging over the potential functions (line 12). For a visual representation of this procedure, please refer to Fig. 8.
Algorithm 3 Nested Combinatorial Sequential Monte Carlo
  Input: $\mathbf{Y}=\{Y_{1},\cdots,Y_{N}\}\in\Omega^{NxM}$, $\theta$
1:  Initialization. $\forall k$, $s_{0}^{k}\leftarrow\perp$, $w_{0}^{k}\leftarrow 1/K$.
2:  for $r=1$ to $R=N-1$ do
3:     for $k=1$ to $K$ do
4:        Resample   $\mathbb{P}(a_{r-1}^{k}=i)=\frac{w_{r-1}^{i}}{\sum_{l=1}^{K}w_{r-1}^{l}}$
5:        for $i=1$ to $L={N-r\choose 2}$ do
6:           for $m=1$ to $M$ do
7:              Form look-ahead partial state
$s_{r}^{k,m}[i]\sim q(\cdot|s_{r-1}^{a_{r-1}^{k}})$
8:              Compute potentials
$w_{r}^{k,m}[i]=\frac{\pi(s_{r}^{k,m}[i])}{\pi(s_{r-1}^{a_{r-1}^{k}})}\cdot\frac{\nu^{-}(s_{r-1}^{a_{r-1}^{k}})}{q(s_{r}^{k,m}[i]|s_{r-1}^{a_{r-1}^{k}})}$
9:           end for
10:        end for
11:        Extend partial state
$s_{r}^{k}$ $=s_{r}^{k,J}[I],$
$\mathbb{P}(I=i,J=j)$ $=\frac{w_{r}^{k,j}[i]}{\sum_{l=1}^{L}\sum_{m=1}^{M}w_{r}^{k,m}[i]}$
12:        Compute weights
$w_{r}^{k}=\frac{1}{ML}\sum_{i=1}^{L}\sum_{m=1}^{M}w_{r}^{k,m}[i]$
13:     end for
14:  end for
  Output: $s_{R}^{1:K}$ , $w_{1:R}^{1:K}$
[…]
Figure 8: Illustration of the Ncsmc framework: In Ncsmc, all possible one-step ahead topologies, which amount to (${N-r\choose 2}$ in total, are systematically enumerated. For the state ${A,B,{C,D}}$, the enumerated topologies include (top): ${A,{B,{C,D}}}$, (center): ${{A,B},{C,D}}$, and (bottom): ${B,{A,{C,D}}}$. Subsequently, $M=1$ sub-branch lengths are stochastically sampled for each edge. Following this, the sub-weights or potentials are computed (right), and a single candidate is randomly selected proportional to its sub-weight (or potential) to create the new partial state.
Appendix D Approximate Posteriors
The proposal distribution for our point estimator adaptation of Csmc and the corresponding approximate posterior for Vcsmc corresponding to Eq. 5 can be written explicitly as follows:
$Q_{\phi}\left(\mathcal{T}_{1:R}^{1:K},a_{1:R-1}^{1:K}\right)\coloneqq\left(\prod\limits_{k=1}^{K}q_{\phi}(\mathcal{T}_{1}^{k})\right)\cdot\prod\limits_{r=2}^{R}\prod\limits_{k=1}^{K}\left[\frac{w_{r-1}^{a_{r-1}^{k}}}{\sum_{l=1}^{K}w_{r-1}^{l}}\cdot q_{\phi}\left(\mathcal{T}_{r}^{k}|\mathcal{T}_{r-1}^{a_{r-1}^{k}}\right)\right].$ (22)
In the above, $a_{r-1}^{k}$ denotes the ancestor index of the resampled random variable and the partial state $s_{r}^{k}=\mathcal{T}_{r}^{k}$ is sampled by proposing forest $\mathcal{T}_{r}^{k}\sim q_{\phi}(\cdot|\mathcal{T}_{r-1}^{a_{r-1}^{k}})$ from a Uniform distribution. Similarly, the proposal distribution for the global posterior defined in Eq. 14 is expressed where $\lambda_{r}^{k}\sim q_{\psi}(\cdot|\lambda_{r-1}^{a_{r-1}^{k}})=\log N(\cdot|\tilde{\mu},\tilde{\Sigma})$:
$Q_{\phi,\psi}\left(\mathcal{T}_{1:R}^{1:K},\Lambda_{1:R}^{1:K},a_{1:R-1}^{1:K}\right)\coloneqq\left(\prod\limits_{k=1}^{K}q_{\phi}(\mathcal{T}_{1}^{k})\cdot q_{\psi}(\lambda_{1}^{k})\right)\cdot\prod\limits_{r=2}^{R}\prod\limits_{k=1}^{K}\left[\frac{w_{r-1}^{a_{r-1}^{k}}}{\sum_{l=1}^{K}w_{r-1}^{l}}\cdot q_{\phi}\left(\mathcal{T}_{r}^{k}|\mathcal{T}_{r-1}^{a_{r-1}^{k}}\right)\cdot q_{\psi}\left(\lambda_{r}^{k}|\lambda_{r-1}^{a_{r-1}^{k}}\right)\right].$ (23)
The Ncsmc method detailed in Algorithm 3 can also be used to form an unbiased and consistent estimator of the log-marginal likelihood $\widehat{\mathcal{Z}}_{NCSMC}$ and a variational objective which we refer to as $\mathcal{L}_{NCSMC}$:
[…]
Appendix E Log Conditional Likelihood $\hat{P}(X|\tau,\lambda)$
Figure 9: Log-conditional likelihood $\hat{P}(X|\tau,\lambda)$ values for Vcsmc (blue) with $K=\{8,16,32,64,128,256\}$ samples and Vncsmc (red) with $K=\{8,16,32,64,128,256\}$ and $M=1$ samples averaged across 5 random seeds. Greater values of $K$ result in a more constrained ELBO and higher log-likelihood values while reducing stochastic gradient noise. Vncsmc with $K\geq 8$ explores higher probability spaces than the likelihood returned by the simulator, as depicted by the green trace for reference. Vncsmc achieves convergence in fewer epochs than Vcsmc and yields higher values, all while maintaining lower stochastic gradient noise. Notably, even Vncsmc with $(K,M)=(8,1)$ in the top-left plot (red) outperforms Vcsmc with K = 256 in the bottom-right plot (blue).
Appendix F Implementation Details
F.1 Invalid Partial States when Coalescing Particles
Physics imposes several constraints on which pairs of particles are impossible to coalesce. We must consider these constraints as we are building trees from leaf to root, coalescing particles (represented as nodes) at every iteration. The following are the conditions
  1. 1.
    $t>0$ for any node.
  2. 2.
    $t_{p}>t_{cut}$ for all inner nodes.
  3. 3.
    $t_{p}>max(t_{l},t_{r})$ for all inner nodes.
Recall that the ELBO is a function of the weight matrix, which is of dimensions $(R,K)$, and contains all weights of the $K$ particles across $R$ iterations. Each entry in the matrix represents the corresponding weight of some partial state.
In Vcsmc and Vncsmc, resampling ensures that we not only extend upon partial states of valid non-zero probability, but we also arrive at $K$ valid trees at the final rank event. We note that both Greedy Search and Beam Search often fail to find any valid trees because they reach a set of partial states where no viable tree can be constructed.